# Energy momentum tensor on and off the light cone

### v1

@@ Abstract
We compute the gravitational form factors $A_{i}$, $D_{i}$ and $\bar{c}_{i}$ of a scalar theory using both the light-cone and covariant perturbation theory at the one-loop level. The light-cone formalism provides a potential approach to access these form factors beyond the perturbative regime. However, unlike the covariant formulation, the Poincaré symmetry on the light cone is not manifest. In this work, we use perturbation theory as a benchmark to extract the gravitational form factors from the light-front energy-momentum tensor. By comparing results on and off the light cone, we identify $T^{++},T^{+a},T^{+-},T^{12}$ as the “good currents” that are properly renormalized and can be used to extract the gravitational form factors.
@@ I Introduction
Facing these difficulties in the non-perturbative calculation, we turn to perturbation theory for guidance. Recently, More et. al. evaluated the EMT of a quark dressed by gluons using leading order (LO) light-cone perturbation theory (LCPT) [29, 30]. Their results can be compared with the recent LO QED calculations in covariant perturbation theory (CPT) [31, 32, 33, 34]. More et. al. suggested to use spin-flip HMEs to extract GFFs $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$, which are unfortunately not available in the scalar case. Furthermore, they obtained a non-vanishing $\bar{c}(Q^{2})$ except at $Q=0$. They attributed the violation of current conservation to zero modes [30]. In this work, we compute the EMT of the scalar Yukawa theory using both covariant and LCPT at the one-loop level. Scalar theory is among the simplest yet non-trivial quantum field theories in 3+1 dimensions [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 19, 47]. The zero modes within the scalar theory can be removed using Pauli-Villars regularization [48, 16]. Therefore, it is an ideal testbed for developing computational techniques in light-front field theory. By comparing the covariant and light-cone perturbation theory for the scalar Yukawa model, we identify the sources of the divergences and explore how to properly extract GFFs $A_{i}(Q^{2})$, $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$ from the light-front EMT.
@@ III.1 Covariant perturbation theory
where, $\Delta_{e}=(x+y)^{2}m^{2}+(1-x-y)\mu^{2}-xyq^{2}$, $\Delta_{f}=(1-x-y)^{2}m^{2}+(x+y)\mu^{2}-xyq^{2}$ and $\mathcal{O}^{\mu\nu}_{e}=g^{\mu\nu}\ell^{2}_{E}+2(1-x-y)^{2}2P^{\mu}P^{\nu}+((x-y)^{2}-1)(q^{\mu}q^{\nu}-q^{2}g^{\mu\nu})+2g^{\mu\nu}\Big{\{}\big{[}1-(1-x-y)^{2}\big{]}m^{2}+\frac{1}{2}(x^{2}-x+y^{2}-y)q^{2}\Big{\}}$, $\mathcal{O}^{\mu\nu}_{f}=\ell^{\mu\nu}_{E}+2(1-x-y)^{2}2P^{\mu}P^{\nu}+\big{[}(x-y)^{2}-1\big{]}(q^{\mu}q^{\nu}-g^{\mu\nu}q^{2})+2g^{\mu\nu}\big{\{}\mu^{2}-(1-x-y)^{2}m^{2}+\frac{1}{2}(x^{2}-x+y^{2}-y)q^{2}\big{\}}$.
@@ III.2 Light-cone perturbation theory
Since individually,
@@
the violation of the current conservation leads to a non-vanishing total $\bar{c}$:
$\bar{c}(-q^{2})\equiv\sum_{i}\bar{c}_{i}(-q^{2})=\frac{Z}{m^{2}}\Sigma(m^{2})\,.$ (43)
Note that $\bar{c}$ is also divergent. Fortunately, we can compare with the covariant results diagram by diagram to pinpoint the sources of the violation. We find that the violation of the current conservation originates from $t^{11}_{i},t^{22}_{i}$ with contributions from diagrams ($e$) and ($f$). The origin of the discrepancy is that the counterterms derived from renormalizing the Hamiltonian operator are not sufficient to remove the divergences in $t^{11},t^{22}$. This is in contrast to $t^{+-}$ which in fact involves the same operator as the Hamiltonian. On the other hand, $t^{+1}_{i},t^{+2}_{i},t^{12}_{i}$ are identical to the respective covariant results. The minus HMEs, $t^{-i}_{i},t^{--}_{i}$ are very singular and cannot be regularized with a single Pauli-Villars pion. More copies of the Pauli-Villars ghosts are needed to tame these components [47].
@@ IV Covariant light-front dynamics
The presence of $\omega$ also adds one more scalar $\zeta=(\omega\cdot q)/(\omega\cdot P)$ into the calculation. Now, the form factors depend on two Lorentz scalars, $q^{2}$ and $\zeta$. It can be shown that Hermiticity of the EMT operator leads to $F_{i}(\zeta,q^{2})=F^{*}_{i}(-\zeta,q^{2})$, where $F_{i}$ is a form factor [28]. In the Drell-Yan frame $\zeta=0$ and the form factors become real functions that depend on $q^{2}$ only. This choice dramatically simplifies the Lorentz decomposition in covariant light-front dynamics [85]. We will adopt the Drell-Yan frame unless otherwise stated. Then, one of the most general Lorentz decompositions of the HMEs of the EMT operator for a spin-0 hadron on the light front in the Drell-Yan frame $q^{+}=0$ is [^1] This decomposition differs from our previous work [19] by replacing tensor $V^{\alpha}V^{\beta}$ with $V^{\alpha}V^{\beta}+q^{\alpha}q^{\beta}$. Another minor difference is to replace ${q^{4}\omega^{\alpha}\omega^{\beta}}/{(\omega\cdot P)^{2}}$ with ${M^{4}\omega^{\alpha}\omega^{\beta}}/{(\omega\cdot P)^{2}}$. The physical difference of these two choices will be analyzed below.,
@@
where, $P=(p^{\prime}+p)/2$, $q=p^{\prime}-p$. Vector $V^{\alpha}$ is defined as $V^{\alpha}=\varepsilon^{\alpha\beta\rho\sigma}P_{\beta}q_{\rho}\omega_{\sigma}/(\omega\cdot P)$. This vector is perpendicular to $q^{\alpha}$, $P^{\alpha}$ as well as $\omega^{\alpha}$. Hermiticity excludes terms like $P^{\{\alpha}q^{\beta\}}$, $\omega^{\{\alpha}q^{\beta\}}$ or $\omega^{\{\alpha}V^{\beta\}}$. Form factors $S_{1i}$ and $S_{2i}$ are allowed by the reduced kinematic symmetries but forbidden by the full Poincaré symmetry. Therefore, they are called the spurious form factors. They are expected to vanish in the continuum limit. Using this decomposition Eq. (49), we can express the HMEs of the currents in terms of the GFFs as,
@@
has to be satisfied. We note that this relation is not automatically satisfied, except for $q^{2}_{\perp}=0$. Therefore, Newton’s third law (16) holds when integrated over space on the light front $x^{+}=0$. For the scalar Yukawa model, $\bar{c}_{i}(Q^{2})$ are constants in one-loop order. Therefore, Newton’s third law (16) is satisfied everywhere in this case.
@@ Acknowledgements
The authors acknowledge fruitful discussions with J. More, C. Mondal, S. Nair, and K. Serafin.

### v2

@@ Abstract
We compute the gravitational form factors $A_{i}$, $D_{i}$ and $\bar{c}_{i}$ of the scalar Yukawa theory using both the light-cone and covariant perturbation theory at the one-loop level. The light-cone formalism provides a potential approach to access these form factors beyond the perturbative regime. However, unlike the covariant formulation, the Poincaré symmetry on the light cone is not manifest. In this work, we use perturbation theory as a benchmark to extract the gravitational form factors from the light-front energy-momentum tensor. By comparing results on and off the light cone, we identify $T^{++},T^{+a},T^{+-},T^{12}$ as the “good currents” that are properly renormalized and can be used to extract the gravitational form factors.
@@ I Introduction
Facing these difficulties in the non-perturbative calculation, we turn to perturbation theory for guidance. Recently, More et. al. evaluated the EMT of a quark dressed by gluons using leading order (LO) light-cone perturbation theory (LCPT) [29, 30]. Their results can be compared with the recent LO QED calculations in covariant perturbation theory (CPT) [31, 32, 33, 34]. More et. al. suggested to use spin-flip HMEs to extract GFFs $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$, which are unfortunately not available in the scalar case. Furthermore, they obtained a non-vanishing $\bar{c}(Q^{2})$ except at $Q=0$. They attributed the violation of current conservation to zero modes [30]. In this work, we compute the EMT of the scalar Yukawa theory using both covariant and LCPT at the one-loop level. Scalar theory is among the simplest yet non-trivial quantum field theories in 3+1 dimensions [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 19, 47]. The zero modes within the scalar theory are only constrained degrees of freedom [48, 16]. Therefore, it is an ideal testbed for developing computational techniques in light-front field theory. By comparing the covariant and light-cone perturbation theory for the scalar Yukawa model, we identify the sources of the divergences and explore how to properly extract GFFs $A_{i}(Q^{2})$, $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$ from the light-front EMT.
@@ III.1 Covariant perturbation theory
where, $\Delta_{e}=(x+y)^{2}m^{2}+(1-x-y)\mu^{2}-xyq^{2}$, $\Delta_{f}=(1-x-y)^{2}m^{2}+(x+y)\mu^{2}-xyq^{2}$ and $\mathcal{O}^{\mu\nu}_{e}=g^{\mu\nu}\ell^{2}_{E}+2(1-x-y)^{2}2P^{\mu}P^{\nu}+((x-y)^{2}-1)(q^{\mu}q^{\nu}-q^{2}g^{\mu\nu})+2g^{\mu\nu}\Big{\{}\big{[}1-(1-x-y)^{2}\big{]}m^{2}+\frac{1}{2}(x^{2}-x+y^{2}-y)q^{2}\Big{\}}$, $\mathcal{O}^{\mu\nu}_{f}=g^{\mu\nu}\ell^{2}_{E}+2(1-x-y)^{2}2P^{\mu}P^{\nu}+\big{[}(x-y)^{2}-1\big{]}(q^{\mu}q^{\nu}-g^{\mu\nu}q^{2})+2g^{\mu\nu}\big{\{}\mu^{2}-(1-x-y)^{2}m^{2}+\frac{1}{2}(x^{2}-x+y^{2}-y)q^{2}\big{\}}$.
@@ III.2 Light-cone perturbation theory
From the standard Lorentz decomposition Eq. (11), since individually,
@@
the violation of the current conservation leads to a divergent total $\bar{c}$:
$\bar{c}(-q^{2})\equiv\sum_{i}\bar{c}_{i}(-q^{2})=\frac{Z}{m^{2}}\Sigma(m^{2}),$ (43)
in the standard formulation (11). Fortunately, we can compare with the covariant results diagram by diagram to pinpoint the sources of the violation. We find that the violation of the current conservation originates from $t^{11}_{i},t^{22}_{i}$ with contributions from diagrams ($e$) and ($f$). The origin of the discrepancy is that the counterterms derived from renormalizing the Hamiltonian operator are not sufficient to remove the divergences in $t^{11},t^{22}$. This is in contrast to $t^{+-}$ which in fact involves the same operator as the Hamiltonian. On the other hand, $t^{+1}_{i},t^{+2}_{i},t^{12}_{i}$ are identical to the respective covariant results. The minus HMEs, $t^{-i}_{i},t^{--}_{i}$ are very singular and cannot be regularized with a single Pauli-Villars pion. More copies of the Pauli-Villars ghosts are needed to tame these components [47].
@@ IV Covariant light-front dynamics
The presence of $\omega$ also adds one more scalar $\zeta=(\omega\cdot q)/(\omega\cdot P)$ into the calculation. Now, the form factors depend on two Lorentz scalars, $q^{2}$ and $\zeta$. It can be shown that Hermiticity of the EMT operator leads to $F_{i}(\zeta,q^{2})=F^{*}_{i}(-\zeta,q^{2})$, where $F_{i}$ is a form factor [28]. In the Drell-Yan frame $\zeta=0$ and the form factors become real functions that depend on $q^{2}$ only. This choice dramatically simplifies the Lorentz decomposition in covariant light-front dynamics [86]. We will adopt the Drell-Yan frame unless otherwise stated. Then, one of the most general Lorentz decompositions of the HMEs of the EMT operator for a spin-0 hadron on the light front in the Drell-Yan frame $q^{+}=0$ is [^1] This decomposition differs from our previous work [19] by replacing tensor $V^{\alpha}V^{\beta}$ with $V^{\alpha}V^{\beta}+q^{\alpha}q^{\beta}$. The latter violates current conservation. Another minor difference is to replace ${q^{4}\omega^{\alpha}\omega^{\beta}}/{(\omega\cdot P)^{2}}$ with ${M^{4}\omega^{\alpha}\omega^{\beta}}/{(\omega\cdot P)^{2}}$. The physical difference of these two choices will be analyzed below.,
@@
where, $P=(p^{\prime}+p)/2$, $q=p^{\prime}-p$. Vector $V^{\alpha}$ is defined as $V^{\alpha}=\varepsilon^{\alpha\beta\rho\sigma}P_{\beta}q_{\rho}\omega_{\sigma}/(\omega\cdot P)$. This vector is perpendicular to $q^{\alpha}$, $P^{\alpha}$ as well as $\omega^{\alpha}$. Hermiticity excludes terms like $P^{\{\alpha}q^{\beta\}}$, $\omega^{\{\alpha}q^{\beta\}}$ or $\omega^{\{\alpha}V^{\beta\}}$. Form factors $S_{1i}$ and $S_{2i}$ are allowed by the reduced kinematic symmetries but forbidden by the full Poincaré symmetry. Therefore, they are called the spurious form factors. They are expected to vanish in the continuum limit. Note that the structure $V^{\alpha}V^{\beta}+q^{\alpha}q^{\beta}$ violates the current conservation.
Using this decomposition Eq. (49), we can express the HMEs of the currents in terms of the GFFs as,
@@
has to be satisfied. We note that this relation is not automatically satisfied, except for $q^{2}_{\perp}=0$. Therefore, Newton’s third law (16) holds when integrated over space on the light front $x^{+}=0$. For the scalar Yukawa model, $\bar{c}_{i}(Q^{2})$ are constants in one-loop order. Therefore, Newton’s third law (16) is satisfied everywhere in this case. N.B. contrasting to the standard decomposition (11), violation of the current conservation in CLFD (49) is caused by a divergent spurious GFF $S_{2}(Q^{2})=\sum_{i}S_{2i}(Q^{2})$, which is contained in $t^{11}$ and $t^{22}$. This analysis is consistent with the LCPT results in III.2.
@@ Acknowledgements
The authors acknowledge fruitful discussions with J. More, C. Mondal, S. Nair, K. Serafin, and G. A. Miller.